OpenStudy (icalibear):

I'm having trouble with solving system of equations by graphing? Fractions are throwing me off? The problem I need to do is: Solve the system of equations by graphing. y = 1/2x + 3 y = -2x - 7 can somebody please help

OpenStudy (anonymous):

for examplee.... If the function is y=1/4, it is a horizontal line that goes through pt (0, 1/4). If the function is y=1/4x+b (b=constant), then 1/4 is the slope, and that's when you use rise over run. First find a point by plugging in an x and then finding the y. Then from that point, go 1 unit up and 4 units across to get to the next point. Draw a line through those two points and voila!

OpenStudy (anonymous):

hope it helps ... now do the same thing with y = 1/2x + 3

OpenStudy (anonymous):

:D

OpenStudy (icalibear):

k, thanks

OpenStudy (icalibear):

Wouldn't this have no solution then.....? :S ahhh im so confused

OpenStudy (anonymous):